OpenStudy (anonymous):

5x+4y+ 13 and 4x-y= -2

OpenStudy (lieutenantgeneral):

5x+4y+ 13 = plane 4x-y= -2 = y = 2(2x + 1)

OpenStudy (anonymous):

what to do